The chair of an academic department needs to form 3 search committees each consisting of 5 tenured full professors. If the department has 18 such faculty, how many choices
does the chair have if he decides that no one can serve on more than one search committee? Note that a set of 5 professors serving on one search committee is distinguishable
from the same professors serving on another search committee
Answer #1

number of choices =N(select 15 professors from 18 and then divide them in 3 search committees of 5 members each)

=(18C15)*(15!/(5!*5!*5!)) =816*756756 =617512896